# SReferences

### v1

@@ 4.2 Identity-Aware Feature Extractor
$P_{\text{e}}^{\text{ind}}\!=\!\frac{1}{N}\sum_{\text{I}\in\mathbb{I}}\,$ $\mathbb{P}(H_{0}|\text{I})\mathbb{P}(C\!=\!1|H_{0},\text{I})$ (2)
+\,+ $\mathbb{P}(H_{1}|\text{I})\mathbb{P}(C\!=\!0|H_{1},\text{I}),$
@@
where $\mathbb{I}$ is the set of all identities. In Sec. 3 of the supplementary document, we showed that the latter identity-conditioning approach is more powerful in reducing classification error. We conducted a performance comparison between two methods through theory-driven simulations, demonstrating that $P_{\text{e}}^{\text{ind}}$ tends to be lower (better) than $P_{\text{e}}^{\text{com}}$. Furthermore, we observed that the gain of $P_{\text{e}}^{\text{ind}}$ over $P_{\text{e}}^{\text{ind}}$ is more significant when the deepfake traces for individuals are more unique and the detection problem is intrinsically more difficult.
@@
Training. To make the feature extraction network identity-aware, we use a neural network such that the earlier layers extract identity-aware features along with other features, and the later layers extract deepfake traces. We use a learned facial representation, trained by Bulat et al. [14] as the starting point of training $B^{{}^{\prime}}({\cdot})$. Their trained network has an architecture of ResNet. For extracting deepfake features, we tune the portion of the network after the “conv4” block.
@@
We reused the model and initial weights from Bulat et al. [14] for the following three reasons. First, having an existing network that lets personal identity pass through makes our task easier to additionally learn the deepfake traces. In comparison, training a network simultaneously for personal identity and deepfake detection would require joint training of two downstream tasks, which is harder. Second, a deeper network trained with unlabelled data is less biased to any specific portion of the dataset [18]. Bulat et al. [14] pretrained the ResNet architecture with ${\sim}10$ million facial images. Consequently, the initial layers of the network are anticipated to learn a robust representation of features, including the identity. The network is also tested over multiple downstream tasks and therefore, it is a good candidate for extracting facial features [44]. Third, according to Newell and Deng [44], there is an advantage in unsupervised pretraining with unlabeled data when the labeled finetuning dataset is small, which aligns with our labeled training dataset comprising 295 videos from 59 celebrities.
@@
Figure 4: Backbone network training for identity-aware deepfake feature extraction. An authentic and deepfake image pair is passed through the teacher and student networks. Teacher network passes down the deepfake trace knowledge to the student network through loss functions $L_{1}$ and $L_{2}$. The loss function $L_{3}$ increases the feature distance between the authentic and the deepfake image.
@@
The training for the backbone network $B^{{}^{\prime}}(\cdot)$ is depicted in Fig. 4. The input is an image pair, consisting of an authentic image and its corresponding deepfake, generated using a deepfake generation tool. The input is passed through a student network $B_{\rm s}$ and a teacher network $B_{\rm t}$ in parallel. The student network is composed of the pretrained facial representation learning backbone [14] and a concatenated task adaptation head for learning specifically the deepfake traces. The layers after the “conv4” block of the pretrained backbone and the task adaptation head are the tunable portions of the student network. We then utilize the EfficientNetAutoAttB4ST [11] as the teacher network to distill the knowledge for learning the deepfake traces. To adapt the deepfake traces based on personal identity, we add a loss function $L_{3}$ that contrasts the learned traces of a deepfake and its corresponding authentic image in addition to the knowledge distillation losses $L_{1}$ and $L_{2}$. Given the authentic facial image of identity $\rm I$, $\rm f_{auth}$, and its corresponding deepfake image $\rm f_{df}$, the loss terms are defined as follows:
@@
$L_{3}$ $=\Big{[}\max\Big{(}0,m\!-\!D\big{(}B_{\rm s}({\rm f_{auth}}),B_{\rm s}({\rm f_{df}})\big{)}\Big{)}\Big{]}^{2},$ (3c)
where $D(\cdot)$ is the Euclidean distance and $m$ is the margin of the hinge loss. The three loss terms are combined as $\alpha(L_{1}+L_{2})+\beta L_{3}$, with hyperparameters $\alpha$ and $\beta$. $L_{1}$ and $L_{2}$ contribute to the knowledge distillation for learning the deepfake traces and $L_{3}$ contributes to learning the deepfake traces according to identity.
4.3 Identity Decoder for Feature Conditioning
Our identity decoder is a single-layer fully-connected neural network that maps the one-hot-encoded index of a public figure to the feature space generated by our feature extractor. We combine the output of the identity decoder with the output of the identity-aware feature extractor that contains the joint information of the deepfake feature and identity. The extra marginal information provided by the identity decoder can have the effect of conditioning the identity-aware feature, in a similar spirit as in the Bayes rule.
4.4 Contrastive Learning
The Siamese network contains two identical subnetworks that process the two inputs parallelly. The subnetworks learn a manifold for each of the inputs adopting contrastive loss that allows a powerful discrimination between the two inputs. In our work, we designed each of the subnetworks as a single-layer neural network that takes as input the features of the corresponding image and outputs a vector of length $50$. We experimentally verified that this length is enough for discriminating the two cases. Let us call the two subnetworks of the Siamese network $S_{n_{1}}$ and $S_{n_{2}}$, where the first one processes the features of f and the second one processes the features of $R(\text{f})$. We used contrastive loss [31] to train the Siamese network as follows:
@@ 4.4 Contrastive Learning
5 Experimental Results
5.1 Dataset Curation
Deepfake detection methods generally perform well in the in-dataset evaluation, but the performance drops significantly in cross-dataset evaluation [17]. To provide reliable measurements for the performance of the deepfake detection method, we report only the cross-dataset evaluation results. As our deepfake detector conditions on the identity, we need two independent datasets containing facial images of the same set of identities. Most of the public deepfake detection datasets such as DFDC [24], DFD [3], Deeper Forensics [33] do not explicitly mention identity information associated with the videos. It is hard to find the same persons in another dataset, which would be necessary to perform the cross-dataset evaluation of individualized deepfake detection. To the best of our knowledge, our work is the first work that evaluates the deepfake detection performance on another dataset for each person separately. We curate a dataset from Celeb-DF [39] and CACD [16]. Our curated dataset contains $32k$ facial images of $45$ public figures sourced from YouTube videos for the train set and from the cross-age facial image dataset of the same public figures for the test dataset. We plan to release the source code and dataset for public use after publication.
@@ 5.1 Dataset Curation
For the training dataset, we use real videos from the Celeb-DF dataset [39], which is a popular deepfake detection dataset of 59 public figures. We sample frames from the videos at $5$ frames per second (fps), and detect faces from the videos using the MTCNN [56] face detection network. For each individual $i$, we have facial images $f^{\text{cdf}}_{i,j,k}$ from the $j$th authentic Celeb-DF video, where $j\in\{1,...,10\}$, $k\in\{1,...,N_{f}\}$, and $N_{f}$ is the number of the frames extracted from the video.
@@
Examining multiple candidate datasets, we narrowed down to the CACD [16] dataset for cross-dataset evaluation. CACD [16] contains cross-age facial images of 2,000 public figures with an overlap of 45 public figures with the Celeb-DF dataset. From the CACD dataset, we have authentic images $f^{\text{cacd}}_{i,j,k}$ for the $i$th identity and $j$th available age group of that identity, where $j\in\{1,...,10\}$, $k\in\{1,...,N_{i}\}$, and $N_{i}$ is the number of the images available for that age group. To generate deepfake faces for the test set, for each person $i$, we choose another identity $m$ from the database of 2,000 persons and then train a Faceswap-GAN [6] model using the facial images $f^{\text{cacd}}_{i,j,k}$ and $f^{\text{cacd}}_{m,j,k}$.
5.2 Experimental Setup
As our proposed deepfake detector is designed to operate only when the identity is provided, we conduct our evaluation on our curated dataset comprising two distinct subsets of facial images with an overlap of $45$ public figures. We use the cropped facial images from the video frames of Celeb-DF [39] as the training set and the images from CACD [16] as the test set.
@@ 5.2 Experimental Setup
Our proposed method had two stages of training. In the first stage, we trained the identity-aware feature extractor. For this training, we resized the facial images to 224-by-224 and used random cropping and random horizontal flipping for image augmentation. As shown in Fig. 4, we used a pair of images for the backbone training. We enforced identical cropping within the same pair consisting of an authentic image and its deepfake. We used SGD optimizer and a learning rate of $10^{-3}$. The contrastive loss margin was $50$ and the values of $\alpha$ and $\beta$ were varied manually within $[\alpha,\beta]\in(0,1)^{2}$. We set $[\alpha,\beta]$ to $[1,0]$ during the initial 1,500 epochs of training and subsequently modified it to $[0,1]$ for the next 1,500 epochs. In the second stage, we trained the Siamese network and the identity decoder. For this training, we used Adam optimizer, and the contrastive loss margin was $2$ and the learning rate was determined by the grid search within the range of $(10^{-6},10^{-5})$.
@@
5.3 Performance Gain
In this subsection, we compare the performance of our proposed method with two state-of-the-art methods. The first baseline considered is the Xception [20] network trained on the FaceForensics++ dataset [48] with deepfake videos generated by four methods including Faceswap [5]. The second one is the EfficientNetAutoAttB4ST [11] network trained on the DFDC dataset [24], a dataset consisting of deepfake videos generated by various popular face-swapping methods, such as Facewap-GAN [6], StyleGAN [34], Faceswap [5], and NTH [55]. The performance of the two baseline approaches on our test dataset is presented in Tab. 1.
Table 1: Comparison of deepfake detection performance of the proposed method leveraging near-idempotence and identity conditioning with the state of the art.
@@ 5.3 Performance Gain
To ensure a fair comparison with our proposed method, we conducted fine-tuning on these two baseline methods using our training dataset. This involved keeping the features frozen and training a classification layer on top of the features until the performance was saturated on the validation dataset. After finetuning, EfficientNetAutoAttB4ST [11] had an AUC mean of $0.92$ across identities with a sample standard deviation of $0.06$.
@@
To evaluate the idea of utilizing idempotency and identity conditioning, we applied the double neural network operation and obtained the features from our trained identity-aware feature extractor. We concatenated those with the features of EfficientNetAutoAttB4ST [11]. Tab. 1 reveals that the proposed method can achieve an AUC mean of $0.94$ across identities, an increase of $0.02$ from Bonettini et al. [11]. The AUC median across identities was $0.958$ with a gain of $0.032$ from the baseline [11]. The 10%-trimmed mean was $0.947$ with a gain of $0.02$. The AUC standard deviation was reduced by $0.01$ or $17\%$ and the AUC interquartile range was reduced by $0.02$ or $29\%$ compared to the baseline [11]. This result demonstrates that idempotency and identity conditioning can improve performance in validity and variation. The detection results on the test dataset for six of the $45$ public figures are shown in Fig. 5.
@@
Figure 5: ROC curves for deepfake detection using the proposed method. Each plot contains results from a public figure and each curve represents a trial of training the network. AUC values are large with small standard deviations, indicating good performance.
@@
The averaged AUC value among all public figures is $0.940$ and the sample standard deviation is $0.05$. We also performed $t$-tests and the proposed method is significantly better than those of the off-the-shelf detectors in terms of AUC. The larger variance of the AUC values of the baseline methods implies that the deepfake detector may perform convincingly for one identity, but it has a greater risk of exhibiting unacceptable performance for others. This makes the baseline methods less attractive for journalists.
5.4 Ablation Studies

### v2

@@ IV-B1 Motivation
$P_{\text{e}}^{\text{ind}}\!=\!\frac{1}{N}\sum_{\text{I}\in\mathbb{I}}$ $\,\mathbb{P}\left(H_{0}\mid\text{I}\right)\,\mathbb{P}\left(C\!=\!1\mid H_{0},\text{I}\right)$ (2)
$+$ $\,\mathbb{P}\left(H_{1}\mid\text{I}\right)\,\mathbb{P}\left(C\!=\!0\mid H_{1},\text{I}\right),$
@@
where $\mathbb{I}$ is the set of all identities. In Appendix A of the supplementary document, we showed that the latter identity-conditioning approach is more powerful in reducing classification error. We conducted a performance comparison between two methods through theory-driven simulations, demonstrating that $P_{\text{e}}^{\text{ind}}$ tends to be lower (better) than $P_{\text{e}}^{\text{com}}$. Furthermore, we observed that the gain of $P_{\text{e}}^{\text{ind}}$ over $P_{\text{e}}^{\text{com}}$ is more significant when the deepfake traces for individuals are more unique, and the detection problem is intrinsically more difficult.
IV-B2 Training
@@
To make the feature extraction network identity-aware, we use a neural network such that the earlier layers extract identity-aware features along with other features, and the later layers extract deepfake traces. We use a learned facial representation, trained by Bulat et al. [50] as the starting point of training $B^{{}^{\prime}}({\cdot})$. Their trained network has an architecture of ResNet. For extracting deepfake features, we tune the portion of the network after the “conv4” block.
@@
We reused the model and initial weights from Bulat et al. [50] for the following three reasons. First, having an existing network that lets personal identity pass through makes our task easier to additionally learn the deepfake traces. In comparison, training a network simultaneously for personal identity and deepfake detection would require joint training of two downstream tasks, which is harder. Second, a deeper network trained with unlabelled data is less biased to any specific portion of the dataset [48]. Bulat et al. [50] pretrained the ResNet architecture with ${\sim}10$ million facial images. Consequently, the initial layers of the network are anticipated to learn a robust representation of features, including the identity. The network is also tested over multiple downstream tasks, and therefore, it is a good candidate for extracting facial features [49]. Third, according to Newell and Deng [49], there is an advantage in unsupervised pretraining with unlabeled data when the labeled finetuning dataset is small, which aligns with our labeled training dataset.
@@
Figure 4: Backbone network training for identity-aware deepfake feature extraction. An authentic and deepfake image pair is passed through the teacher and student networks. The teacher network passes down the deepfake trace knowledge to the student network through loss functions $L_{1}$ and $L_{2}$. The loss function $L_{3}$ increases the feature distance between the authentic and the deepfake image.
@@
The training for the backbone network $B^{{}^{\prime}}(\cdot)$ is depicted in Fig. 4. The input is an image pair consisting of an authentic image and its corresponding deepfake, generated using a deepfake generation tool. The input is passed through a student network $B_{\rm s}$ and a teacher network $B_{\rm t}$ in parallel. The student network is composed of the pretrained facial representation learning backbone [50] and a concatenated task adaptation head for learning the deepfake traces. The layers after the “conv4” block of the pretrained backbone and the task adaptation head are the tunable portions of the student network. We then utilize the EfficientNetAutoAttB4ST [37] as the teacher network to distill the knowledge for learning the deepfake traces. To adapt the deepfake traces based on personal identity, we add a loss function $L_{3}$ that contrasts the learned traces of a deepfake and its corresponding authentic image in addition to the knowledge distillation losses $L_{1}$ and $L_{2}$. Given the authentic facial image of identity $\rm I$, $\rm f_{auth}$, and its corresponding deepfake image $\rm f_{df}$, the loss terms are defined as follows:
@@
$L_{3}$ $=\Big{[}\max\Big{(}0,m_{\text{h}}\!-\!D\big{(}B_{\rm s}({\rm f_{auth}}),B_{\rm s}({\rm f_{df}})\big{)}\Big{)}\Big{]}^{2},$ (3c)
where $D(\cdot)$ is the Euclidean distance and $m_{\text{h}}$ is the margin of the hinge loss. The three loss terms are combined as $\alpha(L_{1}+L_{2})+\beta L_{3}$, with hyperparameters $\alpha$ and $\beta$. $L_{1}$ and $L_{2}$ contribute to the knowledge distillation for learning the deepfake traces, and $L_{3}$ contributes to learning the deepfake traces according to identity.
IV-C Identity Decoder for Feature Conditioning
Our identity decoder is a single-layer fully-connected neural network that maps the one-hot-encoded index of a public figure to the feature space generated by our feature extractor. We combine the output of the identity decoder with the output of the identity-aware feature extractor that contains the joint information of the deepfake feature and identity. The extra marginal information provided by the identity decoder can have the effect of conditioning the identity-aware feature in a similar spirit to the Bayes rule.
IV-D Contrastive Learning
The Siamese network contains two identical subnetworks that process the two inputs parallelly. The subnetworks learn a manifold for each of the inputs, adopting contrastive loss that allows powerful discrimination between the two inputs. In our work, we designed each of the subnetworks as a single-layer neural network that takes as input the features of the corresponding image and outputs a vector of length $50$. We experimentally verified that this length is enough to discriminate between the two cases. Let us call the two subnetworks of the Siamese network $S_{n_{1}}$ and $S_{n_{2}}$, where the first one processes the features of f and the second one processes the features of $R(\text{f})$. We used contrastive loss [57] to train the Siamese network as follows:
@@ IV-D Contrastive Learning
V Experimental Results
One key difference between our proposed method and the existing literature is the use of the near-idempotence property. This section validates the deepfake operation’s near-idempotence property and experimentally demonstrates the performance gains leveraging this property.
V-A Dataset Curation
The deepfake literature encourages cross-dataset evaluations, as they reveal a significant performance drop compared to in-dataset evaluations [59]. To conduct the cross-dataset evaluation with identity conditioning, we will need two separate datasets containing facial images of the same set of identities. However, the identity information is not included in the existing public deepfake detection datasets. For example, DFDC [60], DFD [61], and Deeper Forensics [62] do not explicitly mention identity information associated with the videos. This makes it difficult to find the same persons from another dataset, which would be necessary to perform the cross-dataset evaluation of individualized deepfake detection. To address this challenge, we curated a dataset with identities and a predefined train–test split, where the training and testing subsets are drawn from two different sources. Using our curated dataset, we report only the cross-dataset evaluation results.
@@ V-A Dataset Curation
Our curated dataset contains 32,000 facial images of $45$ public figures sourced from Celeb-DF [63] for the training subset and from the cross-age facial image dataset [64] of the same public figures for the test dataset. We have publicly released the dataset, which contains authentic and deepfake images of public figures, their names, and a predefined train–test split. For the training subset, we use real videos from the Celeb-DF dataset [63], which is a popular deepfake detection dataset of 59 public figures. We sample frames from the videos at $5$ frames per second (fps), and detect faces from the videos using the MTCNN [65] face detection network. For each individual $i$, we have facial images $f^{\text{cdf}}_{i,j,k}$ from the $j$th authentic Celeb-DF video, where $j\in\{1,...,10\}$, $k\in\{1,...,N_{f}\}$, and $N_{f}$ is the number of the frames extracted from the video.
@@
Examining multiple candidate datasets, we narrowed it down to the CACD [64] dataset for cross-dataset evaluation. CACD [64] contains cross-age facial images of 2,000 public figures with an overlap of 45 public figures with Celeb-DF. From CACD, we have authentic images $f^{\text{cacd}}_{i,j,k}$ for the $i$th identity and $j$th available age group of that identity, where $j\in\{1,...,5\}$, $k\in\{1,...,N_{i}\}$, and $N_{i}$ is the number of the images available for that age group. To generate deepfake faces for the testing subset, for each individual $i$, we choose another identity $m$ from the database of 2,000 persons and then generate deepfake images using Faceswap-GAN [1] and DiffSwap [2] with the facial images $f^{\text{cacd}}_{i,j,k}$ and $f^{\text{cacd}}_{m,j,k}$.
V-B Experimental Setup
@@
Our proposed method had two stages of training. In the first stage, we trained the identity-aware feature extractor. For this training, we resized the facial images to 224-by-224 and used random cropping and random horizontal flipping for image augmentation. As shown in Fig. 4, we used a pair of images for the backbone training. We enforced identical cropping within the same pair, which consisted of an authentic image and its deepfake. We used the minibatch SGD optimizer from PyTorch with a learning rate of $10^{-3}$. The hinge loss margin $m_{\text{h}}$ was set to 50. During the first half of the iterations, a specific pair of values for $(\alpha,\beta)$ was used, which was then switched to a different pair for the second half. In each interval, the values were varied within the $[0,1]\times[0,1]$ region with a grid resolution of $0.25\times 0.25$. The best results were obtained by setting $(\alpha,\beta)=(1,0)$ during the initial 1,500 epochs of training, followed by a modification to $(\alpha,\beta)=(0,1)$ for the subsequent 1,500 epochs. In the second stage, we trained the Siamese network and the identity decoder. For this training, we used Adam optimizer, and the contrastive loss margin $m$ was $2$, and the learning rate was determined by the grid search within the range of $[10^{-6},10^{-5}]$ with a step size of $10^{-6}$.
@@
V-C Baseline Algorithms Selection
Our proposed double neural network-based detector is intended to boost a baseline algorithm. When selecting baseline algorithms, we ensured that they provided reliable numerical performance and also functioned as effective feature extractors for integration with our proposed detector. We initially picked nine state-of-the-art deepfake detection methods and evaluated their performance on our curated dataset to assess their suitability. The model weights of the methods were obtained from the respective authors. The detection performance results are summarized in TABLE I, with a scatter-plot visualization presented in Fig. 1(a).
TABLE I: Performance of deepfake detection methods for protection against Faceswap-GAN [1] and DiffSwap [2].
Method (year)    AUC         AUC
Faceswap-GAN [1] DiffSwap [2]
MesoNet (2018) [33]      0.937         0.522
DSP FWA (2019) [40]     0.818         0.657
Capsule (2019) [34]      0.624         0.653
Xception (2019) [35]      0.792         0.641
Face X-ray (2020) [36]      0.794         0.833
EfficientNet (2020) [37]      0.728         0.717
F3-Net (2020) [41]      0.806         0.694
ICT-Ref (2022) [29]      0.595         0.643
LSDA (2024) [38]      0.843         0.726
Among the nine tested methods, we selected Xception [35] and EfficientNet [37] to evaluate the effectiveness of the double neural network operations for deepfake detection. Although these two may not offer the best performance, they serve as powerful feature extractors for image-related tasks. In contrast, Face X-Ray, LSDA, and DSP-FWA are specialized feature extractors designed for detecting specific artifacts, such as blending boundaries or warping patterns. In addition, F3-Net operates in the frequency domain rather than the image domain, whereas our double neural network-based method compares image features in the spatial domain. EfficientNet and Xception are well-known for their powerful spatial feature extraction capabilities across various tasks, making them ideal choices for this study.
The first baseline considered is the Xception [35] network trained on the FaceForensics++ dataset [66] with deepfake videos generated by four methods, including Faceswap [8]. The second one is the EfficientNetAutoAttB4ST [37] network trained on the DFDC dataset [60], a dataset consisting of deepfake videos generated by various popular face-swapping methods, such as Facewap-GAN [1], StyleGAN [67], Faceswap [8], and NTH [68].
V-D Performance Gain
We investigate the performance gains of a deepfake detector empowered with a double-deepfake operation. First, we evaluate the two baseline approaches on our test dataset without applying the double-deepfake technique against Faceswap-GAN [1] and DiffSwap [2], as shown in TABLE II and TABLE III.
TABLE II: Detection performance of the proposed and baseline methods against Faceswap-GAN [1] generated deepfakes.
@@ V-D Performance Gain
TABLE III: Detection performance of the proposed and baseline methods against DiffSwap [2] generated deepfakes.
Method    AUC AUC AUC trimmed
Mean (SD) Median (IQR)  Mean (10%)
Xception [35] 0.641 (0.10) 0.639 (0.17)     0.641
Xception [35] (tuned) 1.000 (0.00) 1.000 (0.00)     1.000
EfficientNet [37] 0.717 (0.10) 0.724 (0.17)     0.718
EfficientNet [37] (tuned) 1.000 (0.00) 1.000 (0.00)     1.000
Proposed 1.000 (0.00) 1.000 (0.00)     1.000
@@
To ensure a fair comparison with our proposed method, we conducted fine-tuning on these two baseline methods using our training dataset. This involved keeping the features frozen and training a classification layer on top of the features until the performance was saturated on the validation dataset. After finetuning, EfficientNetAutoAttB4ST [37] had an AUC mean of $0.920$ across identities with a sample standard deviation of $0.06$ against Faceswap-GAN.
@@
We applied the double neural network operation to evaluate the idea of utilizing idempotency and identity conditioning and obtained the features from our trained identity-aware feature extractor. We concatenated those with the features of EfficientNetAutoAttB4ST [37]. TABLE II reveals that the proposed method can achieve an AUC mean of $0.940$ across identities, an increase of $0.020$ from Bonettini et al. [37]. The AUC median across identities was $0.958$ with a gain of $0.032$ from the baseline [37]. The 10%-trimmed mean was $0.947$ with a gain of $0.02$. The AUC standard deviation was reduced by $0.01$ or $17\%$, and the AUC interquartile range was reduced by $0.02$ or $29\%$ compared to the baseline [37]. This result demonstrates that idempotency and identity conditioning can improve performance in validity and variation. The detection results on the test dataset for six of the $45$ public figures are shown in Fig. 5.
@@
Figure 5: ROC curves for deepfake detection using the proposed method. Each plot contains results from a public figure, and each curve represents a trial of training the network. AUC values are large with small standard deviations, indicating good performance.
@@
The averaged AUC value among all public figures is $0.940$, and the sample standard deviation is $0.05$. We also performed $t$-tests, and the proposed method is significantly better (with $95\%$ confidence interval) than those of the off-the-shelf detectors in terms of AUC. The larger variance of the AUC values of the baseline methods implies that the deepfake detector may perform convincingly for one identity, but it has a greater risk of exhibiting unacceptable performance for others. This makes the baseline methods less attractive for journalism applications.
V-E Ablation Studies